# Adaptive Prompt: Unlocking the Power of Visual Prompt Tuning

### v4

@@ Appendix C Related Work
Mixture of Experts. Building on classical mixture models with adaptive gating mechanism [27, 30, 70], the MoE model has been extensively studied and refined over the years. Notably, subsequent works [11, 59] introduced the MoE layer as an efficient tool for scaling model capacity. Unlike traditional models applying uniform parameters to all inputs, the MoE layer specific parameter subsets for each input, creating a sparsely activated layer that enables significant capacity growth without proportional computational cost [12, 77]. This efficiency and scalability have driven its adoption across diverse domains and tasks [57, 9, 60].
Theory of Mixture of Experts. Although MoEs have been widely used to scale up large models, their theoretical foundations remain under active development. For example, [23] focused on input-free gating Gaussian MoEs and showed that under maximum likelihood estimation, the experts’ convergence rates depend on the algebraic independence of the expert functions. Next, [49, 46] established convergence rates for both density and parameter estimation in Softmax gating Gaussian MoEs, linking these rates to the solvability of polynomial systems under Voronoi-based loss functions. More recently, [47, 48] employed least squares estimation to identify conditions under which expert functions are identifiable. Under these conditions, the resulting estimation rates improve substantially.
@@
Prompting Methods. Prompt tuning [40], initially proposed for language tasks, offers a simpler yet effective alternative by introducing learnable parameters to the input sequence of backbone models, updating only these parameters during fine-tuning [40, 28, 36, 34]. Despite its simplicity, prompt tuning has demonstrated significant performance gains. Recently, visual prompt tuning has emerged as a promising paradigm in PEFT techniques for computer vision. Current advancements in visual prompt tuning focus on engineering improvements, such as minimizing parameter usage [17] and broadening applicability to diverse tasks [71, 61, 72]. However, the theoretical foundations of prompt-based methods remain under-explored. For instance, [19] investigate the relationship between prompt tuning and adapter methods, while [35] analyze these techniques within the context of mixture of experts models. Furthermore, [54] highlight the limitations of prompting, showing that it cannot alter relative attention patterns and instead biases attention layer outputs in a fixed direction.
@@ Appendix D Implementation Details
Table 4: Specifications of the datasets used in our experiments, including 24 datasets from two benchmarks: FGVC and VTAB-1K [75].
Dataset Description # Classes Train Val Test
@@
CUB-200-2011 [66] Fine-grained bird species recognition 200 5,394 600 5,794
NABirds [63] Fine-grained bird species recognition 55 21,536 2,393 24,633
Oxford Flowers [50] Fine-grained flower species recognition 102 1,020 1,020 6,149
Stanford Dogs [31] Fine-grained dog species recognition 120 10,800 1,200 8,580
Stanford Cars [14] Fine-grained car recognition 196 7,329 815 8,041
@@
Datasets specifications. Table 4 provides a summary of the statistics and details of the classification datasets evaluated in the paper. Figure 5 presents image examples from all 24 datasets. Following [28], each FGVC dataset is randomly split into 90% train and 10% val, with the val set used for hyperparameter tuning. For VTAB-1K, we use an 800-200 split for tuning and train on all available data for the final evaluation.
Pre-trained backbones specifications. This study explores two primary groups of Vision Transformer (ViT) models. The first group comprises conventional ViT architectures [8], including ViT-Base, ViT-Large, and ViT-Huge, pre-trained on the large-scale ImageNet-21K [7] dataset. The second group focuses on self-supervised models, specifically Masked Autoencoders (MAE) [20] and Momentum Contrast v3 (MoCo v3) [4], both pre-trained using ImageNet-1K. Unless otherwise specified, we use ViT-B/16 with supervised pre-training on ImageNet-21K by default.
Augmentation. During training, we employ standard image augmentation techniques. For the five FGVC datasets, images are normalized using ImageNet mean and standard deviation, followed by random resized cropping to $224\times 224$ and random horizontal flipping. For the VTAB-1k, images are directly resized to $224\times 224$.
Hyperparameters. We use val set of each dataset to find best prompt length $N_{p}$, kernel size $K$ and hidden dimension $r$ of the feature projector. Following [28], the search range for $N_{p}$ is $\{1,5,10,50,100,200\}$. Given the relatively small resolution of our feature map ${\bm{X}}_{\mathrm{img}}$ (e.g., $14\times 14$ for ViT), the kernel size $K$ is selected from $\{2,3,4\}$. For the hidden dimension $r$, we explore the range $\{4,8,16,32,64,128,256\}$. To identify the optimal learning rate and weight decay, we perform a grid search, consistent with [42, 28]. The learning rate is searched within $\{50,25,10,5,2.5,1,0.5,0.25,0.1,0.05\}$, while the weight decay is chosen from $\{0.01,0.001,0.0001,0.0\}$. We adopt the batch size settings of [28], using 64 and 128. The models are optimized with SGD for 100 epochs, employing a cosine decay learning rate schedule with 10 warmup epochs.
@@ E.1 Per-task Results for VTAB-1K and FGVC
Table 5: VTAB-1K Natural Per-Task Results for ViT-B/16 Supervised Pre-trained on ImageNet-21K. “Number of Wins” in [$\cdot$] denotes comparisons relative to full fine-tuning. “Tuned/Total” is the percentage of parameters tuned for each task. The highest accuracy among all approaches except Full is highlighted in bold.
@@
Table 6: VTAB-1K Specialized Per-Task Results for ViT-B/16 Supervised Pre-trained on ImageNet-21K. “Number of Wins” in [$\cdot$] denotes comparisons relative to full fine-tuning. “Tuned/Total” is the percentage of parameters tuned for each task. The highest accuracy among all approaches except Full is highlighted in bold.
@@
Table 7: VTAB-1K Structured Per-Task Results for ViT-B/16 Supervised Pre-trained on ImageNet-21K. “Number of Wins” in [$\cdot$] denotes comparisons relative to full fine-tuning. “Tuned/Total” is the percentage of parameters tuned for each task. The highest accuracy among all approaches except Full is highlighted in bold.
@@
Table 8: FGVC Per-Task Results for ViT-B/16 Supervised Pre-trained on ImageNet-21K. “Number of Wins” in [$\cdot$] denotes comparisons relative to full fine-tuning. “Tuned/Total” is the percentage of parameters tuned for each task. The highest accuracy among all approaches except Full is highlighted in bold.
@@ E.2 Per-task Results on MAE and MoCo v3
Tables 9, 10, and 11 provide per-task results for MAE [20], as summarized in Table 2. Similarly, Tables 12, 13, and 14 present per-task results for MoCo v3 [4], corresponding to the overview in Table 2.
E.3 Different Backbone Scales
Figure 6: Comparison of VAPT and VPT across different backbone scales.
In Figure 6, we compare VAPT and VPT on the CUB-200-2011 dataset [66] under varying scales of the pre-trained backbone (ViT-Base, ViT-Large, and ViT-Huge). The results demonstrate that VAPT consistently surpasses VPT as the model size increases. Notably, with the ViT-Huge backbone, VAPT achieves up to a 1.3% improvement over VPT. These findings highlights the scalability and effectiveness of VAPT compared to VPT as model size grows.
E.4 Ablation Study
@@
Impact of different components. We conducted ablation studies on the VTAB-1K benchmark [75] to evaluate the individual contributions of each component in VAPT. The results, presented in Table 15, show that incorporating the channel-wise convolution layer increases the overall average performance by 0.97%. This improvement highlights the benefit of explicitly encoding spatial relationships in the feature map before it is processed by the token-wise projectors. Furthermore, the channel-wise convolution layer reduces the overall parameter count by downsampling the feature map from $H\times W$ to $H^{\prime}\times W^{\prime}$, where $H^{\prime}=H-K+1$ and $W^{\prime}=W-K+1$. Consequently, the dimensionality for each token-wise projector is also reduced, leading to fewer parameters. When the feature projector is removed, performance decreases; for instance, in VTAB-1K Natural, the performance drops from 81.43% to 79.43%. This highlights the importance of the feature projector. Furthermore, our sharing mechanism for the feature projector not only reduces the number of parameters but also facilitates knowledge transfer between layers, leading to improved performance. Overall, combining all components yielded the best performance, with an average accuracy of 72.91% on VTAB-1K.
@@
Detailed analysis of the channel-wise convolution layer. We evaluated the effectiveness of our channel-wise convolution layer by comparing its performance with that of a standard convolution layer. We also explored alternative strategies for modeling spatial relationships, including an average pooling layer [37]. Notably, average pooling can be considered a special case of our channel-wise convolution layer, where the kernel weights are fixed rather than learned. The comparative results are presented in Table 16. Our channel-wise convolution layer not only reduces the number of parameters relative to a standard convolution layer but also mitigates overfitting, thereby improving overall performance. Furthermore, it surpasses the average pooling layer; for instance, on VTAB-1K Structured, we achieve a 0.93% performance gain. This improvement is attributed to the flexibility of our channel-wise convolution layer compared to average pooling.
@@
Robustness to different hyperparameters. We systematically evaluate the influence of key hyperparameters, including prompt length $N_{p}$, kernel size $K$, and hidden dimension $r$, by conducting experiments on three representative VTAB-1K tasks: Sun397 (Natural), Retinopathy (Specialized), and Clevr/distance (Structured). The results are shown in Figure 7, Figure 8, and Figure 9. As depicted, the optimal hyperparameters vary across tasks. Nevertheless, VAPT consistently achieves higher performance than VPT across a range of hyperparameters. Notably, even with only a single prompt, VAPT maintains its advantage over VPT.
@@
Linear activation in the feature projector. As shown in Appendix B, VAPT preserves its optimal sample efficiency even when the activation function $\sigma$ in the feature projector (see equation (11)) is replaced by a linear identity function. To substantiate this theoretical result, we report the corresponding performance in Table 17, where $\sigma$ is removed. The results demonstrates that the linear version of VAPT remains competitive with the non-linear variant. Notably, it still outperforms VPT by a substantial margin (e.g., 4.30% on the VTAB-1K Structured task), confirming both the theoretical and empirical robustness of our approach.
E.5 Computational Cost
@@
One primary concern when designing prompts as adaptive functions of the input is the associated computational overhead. Although VAPT outperforms VPT, one notable advantage of VPT is its simplicity, as its prompts remain fixed regardless of the input. To examine the trade-offs, we compare both the performance and computational cost of the two methods. Table 18 reports these costs measured in FLOPs (GFLOPs) and MACs (GMACs) for VAPT and VPT across five FGVC datasets. VAPT incurs only a slight increase in computational demands, with FLOPs rising by 0.18% for Stanford Cars and between 0.06% and 0.59% across other datasets. A similar pattern holds for MACs. These modest increases are outweighed by the significant performance gains, and it is noteworthy that VAPT also employs fewer parameters than VPT. Overall, these findings highlight the efficiency of VAPT, delivering robust performance improvements at a minimal additional computational cost.
E.6 Interpretative Visualizations
To facilitate a deeper understanding of our method, we provide visualization inspections to support the advantages of VAPT. Specifically, we use GradCAM [58] to generate attention maps by computing the gradients of a target concept with respect to the model’s final layer. Figure 10 presents examples from five VTAB-1K datasets, namely Sun397, SVHN, Resisc45, Clevr/count, and KITTI/distance, comparing heatmaps generated by VAPT and VPT. These visualizations enable us to examine which regions of the input data each technique prompts the model to focus on. We observe that VAPT can localize relevant image regions more accurately than VPT. For instance, in the Sun397 dataset, while VPT struggles to capture the complete structure of an object, VAPT succeeds in identifying and highlighting its key features. This enhanced localization indicates VAPT’s stronger ability to capture salient visual patterns. Consequently, VAPT not only improves upon VPT’s performance but also enhances the interpretability of the model by providing more coherent and precise visual explanations.

### v5

@@ Appendix C Related Work
Mixture of Experts. Building on classical mixture models with adaptive gating mechanisms [30, 33, 77], the MoE model has been extensively studied and refined over the years. Notably, subsequent works [12, 65] introduced the MoE layer as an efficient tool for scaling model capacity. Unlike traditional models that apply uniform parameters to all inputs, the MoE layer applies specific parameter subsets for each input, creating a sparsely activated layer that enables significant capacity growth without proportional increase in computational cost [13, 86]. This efficiency and scalability have driven its adoption across diverse domains and tasks [63, 10, 66].
Theory of Mixture of Experts. Although MoEs have been widely used to scale up large models, their theoretical foundations remain under active development. For example, [24] focused on input-free gating Gaussian MoEs and showed that under maximum likelihood estimation, the experts’ convergence rates depend on the algebraic independence of the expert functions. Next, [53, 50] established convergence rates for both density and parameter estimation in Softmax-gating Gaussian MoEs, linking these rates to the solvability of polynomial systems under Voronoi-based loss functions. More recently, [51, 52] employed least-squares estimation to identify conditions under which expert functions are identifiable. Under these conditions, the resulting estimation rates improve substantially.
@@
Prompting Methods. Prompt tuning [44], initially proposed for language tasks, offers a simpler yet effective alternative by introducing learnable parameters to the input sequence of backbone models, updating only these parameters during fine-tuning [44, 31, 40, 38]. Despite its simplicity, prompt tuning has demonstrated significant performance gains. Recently, visual prompt tuning has emerged as a promising paradigm in PEFT techniques for computer vision. Current advancements in visual prompt tuning focus on engineering improvements, such as minimizing parameter usage [18] and broadening applicability to diverse tasks [78, 67, 79]. However, the theoretical foundations of prompt-based methods remain underexplored. For instance, [20] investigated the relationship between prompt tuning and adapter methods, while [39] analyzed these techniques within the context of MoE models. Furthermore, [60] highlighted the limitations of prompting, showing that it cannot alter relative attention patterns and instead biases attention layer outputs in a fixed direction.
Adaptive Prompting. Prior studies have explored different adaptive prompting strategies. For example, [28, 35] integrated VPT with meta-learning by clustering the downstream dataset and assigning cluster-specific prompts based on the initial ViT layer’s input ${\bm{X}}^{(0)}$. Our approach, VAPT, distinguishes itself by generating prompts at each layer based on the layer-specific input ${\bm{X}}^{(l)}$. This layer-wise conditioning aligns with our theoretical framework linking MoE and VPT, which dictates dependency on ${\bm{X}}^{(l)}$ for both experts and gating functions, facilitating rigorous analysis. Crucially, VAPT’s motivation stems from the functional properties of prompt experts and their disparity with pre-trained experts, not merely from the aim of using different prompts per input. In contrast, [54, 59, 27], add adaptive components directly to layer inputs. From our MoE perspective, this risks altering the pre-trained experts $f_{1},\dots,f_{N}$ and disrupting their learned knowledge from the pre-training process. Furthermore, this approach fixes the number of adaptive components to that of the pre-trained experts and lacks explicit theoretical grounding. VAPT, conversely, introduces new prompt experts alongside frozen pre-trained ones similar to VPT, preserving knowledge, offering flexibility in the number of experts, and is grounded on a solid theoretical foundation derived from the MoE viewpoint. Finally, while some adaptive prompting methods employ complex image-to-prompt transformations that can obscure analysis of their effectiveness, VAPT utilizes a simple yet effective formulation, enabling more straightforward and rigorous theoretical examination.
@@ Appendix D Implementation Details
Table 4: Specifications of the datasets used in our experiments, which include 24 datasets from two benchmarks: FGVC and VTAB-1K [82].
Dataset Description # Classes Train Validation Test
@@
CUB-200-2011 [72] Fine-grained recognition of bird species 200 5,394 600 5,794
NABirds [69] Fine-grained recognition of bird species 55 21,536 2,393 24,633
Oxford Flowers [55] Fine-grained recognition of flower species 102 1,020 1,020 6,149
Stanford Dogs [34] Fine-grained recognition of dog species 120 10,800 1,200 8,580
Stanford Cars [15] Fine-grained recognition of car 196 7,329 815 8,041
@@
Dataset Specifications. Table 4 provides a summary of the statistics and details of the classification datasets evaluated in this paper. Figure 3 presents image examples from all 24 datasets. Following [31], each FGVC dataset is randomly split into 90% train and 10% val, with the val set used for hyperparameter tuning. For VTAB-1K, we use an 800-200 split for tuning, and we train on all available data for the final evaluation.
Pre-trained Backbone Specifications. This study explores two primary groups of Vision Transformer (ViT) models. The first group comprises conventional ViT architectures [9], including ViT-Base, ViT-Large, and ViT-Huge, pre-trained on the large-scale ImageNet-21K [7] dataset. The second group focuses on self-supervised models, specifically Masked Autoencoders (MAE) [21] and Momentum Contrast v3 (MoCo v3) [4], both pre-trained on ImageNet-1K. Unless otherwise specified, we use ViT-B/16 with supervised pre-training on ImageNet-21K dataset by default.
Augmentation. During training, we employ standard image augmentation techniques. For the five FGVC datasets, images are normalized using ImageNet mean and standard deviation, followed by random resized cropping to $224\times 224$ and random horizontal flipping. For VTAB-1k, images are directly resized to $224\times 224$.
Hyperparameters. We use the val set of each dataset to find the optimal prompt length $N_{p}$, kernel size $K$, and hidden dimension $r$ of the feature projector. Following [31], the search range for $N_{p}$ is $\{1,5,10,50,100,200\}$. Given the relatively small resolution of our feature map ${\bm{X}}_{\mathrm{img}}$ (e.g., $14\times 14$ for ViT), the kernel size $K$ is selected from $\{2,3,4\}$. For the hidden dimension $r$, we explore the range $\{4,8,16,32,64,128,256\}$. To identify the optimal learning rate and weight decay, we perform a grid search, consistent with [46, 31]. The learning rate is searched within $\{50,25,10,5,2.5,1,0.5,0.25,0.1,0.05\}$, while the weight decay is chosen from $\{0.01,0.001,0.0001,0.0\}$. We adopt the batch size settings of [31], using values of 64 and 128. The models are optimized with SGD for 100 epochs, employing a cosine decay learning rate schedule with 10 warm-up epochs.
@@ E.1 Per-task Results for VTAB-1K and FGVC
Table 5: VTAB-1K Natural per-task results for ViT-B/16 supervised pre-trained on ImageNet-21K. “Number of Wins” in [$\cdot$] denotes the number of wins relative to full fine-tuning. “Tuned/Total” represents the percentage of parameters tuned for each task. The highest accuracy among all approaches except Full is highlighted in bold.
@@
Table 6: VTAB-1K Specialized per-task results for ViT-B/16 supervised pre-trained on ImageNet-21K. “Number of Wins” in [$\cdot$] denotes the number of wins relative to full fine-tuning. “Tuned/Total” represents the percentage of parameters tuned for each task. The highest accuracy among all approaches except Full is highlighted in bold.
@@
Table 7: VTAB-1K Structured per-task results for ViT-B/16 supervised pre-trained on ImageNet-21K. “Number of Wins” in [$\cdot$] denotes the number of wins relative to full fine-tuning. “Tuned/Total” represents the percentage of parameters tuned for each task. The highest accuracy among all approaches except Full is highlighted in bold.
@@
Table 8: FGVC per-task results for ViT-B/16 supervised pre-trained on ImageNet-21K. “Number of Wins” in [$\cdot$] denotes the number of wins relative to full fine-tuning. “Tuned/Total” represents the percentage of parameters tuned for each task. The highest accuracy among all approaches except Full is highlighted in bold.
@@ E.2 Per-task Results on MAE and MoCo v3
Tables 9, 10, and 11 provide per-task results for MAE [21], as summarized in Table 2. Similarly, Tables 12, 13, and 14 present per-task results for MoCo v3 [4], which are also summarized in Table 2.
E.3 Statistical Significance Tests
Table 15: Wilcoxon signed-rank test evaluating whether VAPT significantly outperforms other methods across 19 VTAB tasks. The results show that VAPT is statistically significantly better than other baselines ($p<0.05$).
Method Full Bias Adapter VPT E2VPT
$p$-value 5.7e-06 1.9e-06 1.9e-06 1.9e-06 0.0001
To determine whether VAPT consistently surpasses competing PEFT methods across the 19 VTAB tasks, we conducted a one-tailed paired Wilcoxon signed-rank test [74] for each VAPT-baseline comparison. The null hypothesis $H_{0}$ for each comparison was that the median difference in performance scores between VAPT and the baseline method is zero. The alternative hypothesis $H_{1}$ was that VAPT performs better than the baseline method. As reported in Table 15, all resulting $p$-values were below the 0.05 significance level, indicating that the observed performance improvements of VAPT are statistically significant.
E.4 Different Backbone Scales
Figure 4: Comparison of VAPT and VPT on CUB-200-2011 across different backbone scales.
In Figure 4, we compare VAPT and VPT on the CUB-200-2011 dataset [72] using pre-trained backbones of varying scales (ViT-Base, ViT-Large, and ViT-Huge). The results demonstrate that VAPT consistently surpasses VPT as the model size increases. Notably, with the ViT-Huge backbone, VAPT achieves up to a 1.3% improvement over VPT. These findings highlight the scalability and effectiveness of VAPT compared to VPT as model size grows.
E.5 Semantic Segmentation
Table 16: Semantic segmentation results on ADE20K. All methods are evaluated with SETR [84] using ViT-L. The best mIoU scores among all methods but Full are bolded.
Method Full Head Bias VPT VAPT
mIoU 48.31 35.12 43.40 42.24 44.04
# params 318.31 13.18 13.46 15.60 15.29
To explore the potential generalizability of VAPT beyond visual classification, we evaluated its performance on the semantic segmentation task. We report mIoU values in Table 16. Our observations indicate that VAPT attains a higher mIoU than VPT while introducing fewer additional parameters. Furthermore, VAPT remains competitive with other PEFT approaches. These findings highlight VAPT’s strong generalizability across various computer vision tasks.
E.6 Ablation Study
@@
Impact of Different Components. We conducted ablation studies on the VTAB-1K benchmark [82] to evaluate the individual contributions of each component in VAPT. The results, presented in Table 17, show that incorporating the channel-wise convolution layer increases the overall average performance by 0.97%. This improvement highlights the benefit of explicitly encoding spatial relationships in the feature map before it is processed by the token-wise projectors. Furthermore, the channel-wise convolution layer reduces the overall parameter count by downsampling the feature map from $H\times W$ to $H^{\prime}\times W^{\prime}$, where $H^{\prime}=H-K+1$ and $W^{\prime}=W-K+1$. Consequently, the dimensionality for each token-wise projector is also reduced, leading to fewer parameters. When the feature projector is removed, performance decreases; for instance, in VTAB-1K Natural, the performance drops from 81.43% to 79.43%. This decrease highlights the importance of the feature projector. Furthermore, our sharing mechanism for the feature projector not only reduces the number of parameters but also facilitates knowledge transfer between layers, leading to improved performance. Overall, combining all components yielded the best performance, with an average accuracy of 72.91% on VTAB-1K.
@@
Detailed Analysis of the Channel-wise Convolution Layer. We evaluated the effectiveness of our channel-wise convolution layer by comparing its performance to that of a standard convolution layer. We also explored alternative strategies for modeling spatial relationships, including an average pooling layer [41]. Notably, average pooling can be considered a special case of our channel-wise convolution layer, where the kernel weights are fixed rather than learned. The comparative results are presented in Table 18. Our channel-wise convolution layer not only reduces the number of parameters compared to a standard convolution layer but also mitigates overfitting, thereby improving overall performance. Furthermore, it surpasses the average pooling layer; for instance, on VTAB-1K Structured, we achieve a 0.93% performance gain. This improvement is attributed to the greater flexibility of our channel-wise convolution layer compared to that of average pooling.
@@
Robustness to Different Hyperparameters. We systematically evaluated the influence of key hyperparameters, including prompt length $N_{p}$, kernel size $K$, and hidden dimension $r$, by conducting experiments on three representative VTAB-1K tasks: Sun397 (Natural), Retinopathy (Specialized), and Clevr/distance (Structured). The results are shown in Figures 56, and 7. As depicted, the optimal hyperparameters vary across tasks. Nevertheless, VAPT consistently achieves higher performance than VPT across a range of these hyperparameters. Notably, even with only a single prompt, VAPT maintains its advantage over VPT.
@@
Linear Activation in the Feature Projector. As shown in Appendix B, VAPT preserves its optimal sample efficiency even when the activation function $\sigma$ in the feature projector (see Equation (11)) is replaced by a linear identity function. To substantiate this theoretical result, we report the corresponding performance in Table 19, where $\sigma$ is removed. The results demonstrate that the linear version of VAPT remains competitive with the non-linear variant. Notably, it still outperforms VPT by a substantial margin (e.g., 4.30% on the VTAB-1K Structured task), confirming both the theoretical and empirical robustness of our approach.
E.7 Computational Cost
@@
One primary concern when designing prompts that adapt to the input is the associated computational overhead. Although VAPT outperforms VPT, one notable advantage of VPT is its simplicity, as its prompts remain fixed regardless of the input. To examine these trade-offs, we compare both the performance and computational costs of the two methods. Table 20 reports these costs, measured in FLOPs (GFLOPs) and MACs (GMACs) for VAPT and VPT across five FGVC datasets. VAPT incurs only a slight increase in computational cost, with FLOPs increasing by 0.18% for Stanford Cars and ranging between 0.06% and 0.59% across other datasets. A similar pattern is observed for MACs. These modest increases are outweighed by the significant performance gains. Notably, VAPT also employs fewer parameters than VPT. Overall, these findings highlight the efficiency of VAPT, delivering robust performance improvements at a minimal additional computational cost.
E.8 Interpretive Visualizations
To facilitate a deeper understanding of our method, we provide visualizations to illustrate the advantages of VAPT. Specifically, we use GradCAM [64] to generate attention maps by computing the gradients of a target concept with respect to the model’s final layer. Figure 8 presents examples from five VTAB-1K datasets, namely Sun397, SVHN, Resisc45, Clevr/count, and KITTI/distance, comparing heatmaps generated by VAPT and VPT. These visualizations enable us to examine the regions of the input data to which each technique directs the model’s focus. We observe that VAPT can localize relevant image regions more accurately than VPT. For instance, in the Sun397 dataset, while VPT struggles to capture the complete structure of an object, VAPT succeeds in identifying and highlighting its key features. This enhanced localization indicates VAPT’s stronger ability to capture salient visual patterns. Consequently, VAPT not only improves performance over VPT but also enhances the interpretability of the model by providing more coherent and precise visual explanations.